Sam needs to get carpet for two rooms of his house. Estimate the number of square feet (square footage) in the two rooms if one room measures 10 1/2 feet by 9 3/4 feet; and the other room measures 19 1/4 feet by 18 1/2 feet.
Answer by checkley71(8403)

10.5*9.75=102.375 SQ FEET.
19.25*18.5=356.125 SQ FEET.
102.375+356.125=458.5 TOTAL SQ FEET OF CARPET IS NEEDED.